# Geometric Autoencoder Priors for Bayesian Inversion: Learn First Observe Later

### v2

@@ Abstract
Uncertainty Quantification (UQ) is paramount for inference in engineering. A common inference task is to recover full-field information of physical systems from a small number of noisy observations, a usually highly ill-posed problem. Sharing information from multiple distinct yet related physical systems can alleviate this ill-possendess. Critically, engineering systems often have complicated variable geometries prohibiting the use of standard multi-system Bayesian UQ. In this work, we introduce Geometric Autoencoders for Bayesian Inversion (GABI), a framework for learning geometry-aware generative models of physical responses that serve as highly informative geometry-conditioned priors for Bayesian inversion. Following a “learn first, observe later” paradigm, GABI distills information from large datasets of systems with varying geometries, without requiring knowledge of governing PDEs, boundary conditions, or observation processes, into a rich latent prior. At inference time, this prior is seamlessly combined with the likelihood of a specific observation process, yielding a geometry-adapted posterior distribution. Our proposed framework is architecture agnostic. A creative use of Approximate Bayesian Computation (ABC) sampling yields an efficient implementation that utilizes modern GPU hardware. We test our method on: steady-state heat over rectangular domains; Reynold-Averaged Navier-Stokes (RANS) flow around airfoils; Helmholtz resonance and source localization on 3D car bodies; RANS airflow over terrain. We find: the predictive accuracy to be comparable to deterministic supervised learning approaches in the restricted setting where supervised learning is applicable; UQ to be well calibrated and robust on challenging problems with complex geometries.
@@ 1 Introduction
In this work, we propose a methodology using graph-based autoencoders to learn highly informative priors as geometry-aware generative models for use in Bayesian inversion tasks. We also show how to effectively solve the Bayesian inverse problems resulting from these priors through a pushfoward model. Importantly, the proposed scheme decouples the task of “learning” and the task of ”inferring from observations”, resulting in train-once use everywhere approach which is independent of the particular observation process for solving inverse problems, a trait not shared by related supervised learning approaches. We make the following contributions:
@@
    We demonstrate the scalability of the method on a large terrain flow problem through a multi-GPU implementation; showing potential for training GABI foundation models.
@@ 1.1 Setup
  • In the application stage, we use this learned prior to solve inverse problems for data coming from new physical systems not in the dataset $\mathcal{D}$. That is, given sparse and noisy observations $\mathbf{y}_{o}$ of a solution $\mathbf{u}_{o}$ of a physical systems with geometry $\mathsf{M}_{o}$, we use our previously learned prior to solve the Bayesian inverse problem for the unknown full-field $\mathbf{u}_{o}$. We recover a distribution over the solution $\mathbf{u}_{o}$ by only observing $\mathbf{y}_{o}$.
Using the proposed methodology, one can encode prior beliefs through geometry-dependent full-field solution datasets. Practitioners can then construct highly-informative priors from data across varying physical geometries. This prior is independent of the observation process for any particular inference task, hence this is a learn-once apply everywhere approach.
@@ 1.2.2 Forward and Inverse Problems on Graphs
Graphs are handled in machine learning frameworks through geometric deep learning (Bronstein et al., 2021; Wu et al., 2022; Kipf, 2016). Several important works consider inverse problems where the quantity inferred lives of a graph description of the geometry. In Garcia Trillos & Sanz-Alonso (2018) the authors prove the convergence of the graph posterior to the continuum posterior subject to some conditions. In Povala et al. (2022) the adjacency structure of a mesh is leveraged to accelerate variational inference inversion. Important works consider solving the PDE forward problem through physics informed graph machine learning(Seo & Liu, 2019; Valencia et al., 2025; Wandel et al., 2020) Used in our comparisons are linear inverse problems with Gaussian process priors, this leads to graph Gaussian process methods (Borovitskiy et al., 2021; Mostowsky et al., 2024). We find it important to mention works in the physics informed literature where (assuming complete or partial knowledge on PDEs, boundary conditions, forcing etc.) one can use physics residuals to learn forward and inverse maps for PDEs (Raissi et al., 2019; Gao et al., 2022; Vadeboncoeur et al., 2023; Xu et al., 2023; Rixner & Koutsourelakis, 2021). However, in this work, we focus solely on methods which do not assume knowledge of the governing PDE, boundary conditions, or forcing and are entirely data-driven in their approach.
@@ 1.3 Notation
We denote by $\mathbb{P}_{z}$ a probability measure on the measurable space $(\mathcal{Z},\mathcal{S})$, and by $p_{z}$ the associated density with respect to Lebesgue measure when it exists. Empirical measures are a mixture of Dirac measures. That is, for $D=\{x_{i}\}_{i=1}^{M}$, we have $\widehat{\mathbb{E}}_{x\in D}[\delta_{x}]=\frac{1}{M}\sum_{i=1}^{M}\delta_{x_{i}}$ where $\delta_{x_{i}}(A)=1\text{ if }x_{i}\in A,\text{ and }\delta_{x_{i}}(A)=0,\text{ if }x_{i}\notin A.$ We will sometimes refer to such probability measures as the “densities” $p,q$ to lighten notation. Superscripts in parentheses denote realizations of random variables. For a function $g:\mathcal{Z}\rightarrow\mathcal{U}$, the pushforward $g_{\#}p_{z}$ means $g_{\#}\mathbb{P}_{z}(A)=\mathbb{P}_{z}(g^{-1}(A))$ for all measurable sets $A$. We note, a sample $u^{(n)}\sim g_{\#}p_{z}$ is obtained as $u^{(i)}=g(z^{(i)}),z^{(i)}\sim p_{z}$.
@@ 2.1 Posterior from a Pushforward Prior
Stated simply, in Bayesian inference, when a prior on $\mathcal{U}$ is specified as the pushforward under $g$ (Subsection 1.3) of a latent prior on a space $\mathcal{Z}$, two distributions are equivalent: i) the Bayesian posterior on $\mathcal{U}$, and ii) the pushforward under $g$ of the posterior associated to the latent prior on $\mathcal{Z}$. This result will be explicitly related to the geometric inference setting in Theorem 2.2, showing that with a latent prior constructed via a geometry aware encoder/decoder, we can solve inverse problems on $\mathcal{U}$ with a single common latent representation relevant to any number of in-distribution geometries. All proofs can be found in Appendix A.
@@ 3 Implementation
Neural Networks: We make use of meshes and graph neural networks to describe and manipulate geometric information. The proposed methodology is agnostic to any particular architecture, however we need the following characteristics: Geometry aware architecture and data-structure; Mapping to and from fixed-dimensional vectors; Non-locality. We expand on these points in B. For the heat, airfoil, and car problems we make use of a Graph Convolutional Network with interleaved nonlocal averaging layers (Lanthaler et al., 2024). For the terrain flow problem we use a GEN (Li et al., 2020) GNN architecture. We detail these architectures in Appendix B.1. We also test a Transformer on the heat problem in Appendix B.3.3.
@@ 4.2 Airfoils – Flow Field
In this section, we test the proposed methodology on the full field reconstruction or airflow around an airfoil from a dataset of Reynolds averaged Navier-Stokes simulations. We use the setup described in Duthé et al. (2025). Here, the geometric autoencoder is trained on the pressure and 2D velocity field of 1k airfoil geometries with various inflow conditions. Once trained, the observation vector $\mathbf{y}_{o}$ is taken to be noisy pressure observations at a small number of nodes on the surface of the airfoil, emulating a physical scenario when such pressure sensors may be attached to a wing/blade. The Bayesian inversion task is to reconstruct the pressure and velocity field around the airfoil, a task far too ill-posed without a highly informative prior. To compile the results in Table 3 we test GABI-ABC against the Direct Map with the same architecture type – we randomize the number of observation locations to be drawn between 5 and 50 observations with probability inversely proportional to the number of observation locations. For the supervised Direct Map approach the distribution of the number of observation must be known at train time, this is not the case for GABI who’s training is independent of the observation process. Figure 3 shows a select number of results. For a discussion on out-of-distribution inference we refer the reader to Appendix B.4.2.
@@ 4.3 Car Body – Acoustic Vibration and Source Localization
Figure 4: Ground truth, inferred mean, stddev., and error for ampltitude $u$, and forcing $f$. In magenta are the observation locations.
@@ 5 Conclusion
We propose a methodology for learning highly informative priors for solving Bayesian inverse problems across varying geometries. The scheme is realized through a geometric autoencoder which learns informative latent representations allowing for sharing of information across geometries. Once trained this foundation model can be used with any observation process to perform Bayesian full field reconstruction over new geometries. A creative use of ABC sampling leads to an efficient GPU implementation. The method achieves similar predictive accuracy to supervised direct map methods, all while outputting a full Bayesian posterior over the fields of interest and being far more flexible due to the model being independent from the observation process. The method is tested on a wide range of physical setups and geometries.
@@ Appendix A Proofs
This follows directly from Lemma 2.1 and Theorem 2.2 by identifying $z$ with $(\mathbf{z},\sigma_{o})$, $\mathbb{P}_{z}$ with $(\mathbb{P}_{\mathbf{z}}\otimes\mathbb{P}_{\sigma_{o}})$, $g$ with $(D^{\psi}_{o}\otimes\mathrm{Id}.)$, changing to potential function to $\Phi(\mathbf{u}_{o},\sigma_{o};\mathbf{y}_{o})=\frac{1}{2\sigma_{o}^{2}}\|\mathbf{y}_{o}-\mathbf{H}_{o}\bf{\mathbf{u}}_{o}\|^{2}_{2}$, and finally noting $\mathrm{Id}._{\#}\mathbb{P}_{\sigma_{o}}=\mathbb{P}_{\sigma_{o}}$. ∎
@@ Appendix B Additional Numerical Results and Details
The proposed methodology is architecture agnostic, however for an architecture to be appropriate for the task we require a few particular qualities:
  • Geometry aware architecture and data-structure. As the main premise of this work is to share information about physical fields which depend on the geometry of the problem to which they relate, it is essential that the data-structure reflect this geometric information, that the neural network architecture be effective at processing it, and that both naturally accommodate inputs of varying size and topology.
  • Mapping to and from fixed-dimensional vectors. This framework relies on learning latent representations between solution to PDEs and the geometries on which these are defined. Hence, one needs to map variable-sized geometric inputs to and from fixed dimensional latent spaces.
@@ B.1 Graph Convolutional Network
We use graph convolutional layers (Kipf, 2016) in the examples heat on rectangles, RANS around airfoil, Helmholtz on car. To induce non-locality of mappings across channels, $c$, at each layer we set $\texttt{GCN}_{l}:\mathsf{M}_{n}\times\mathbb{R}^{2c}\rightarrow\mathsf{M}_{n}\times\mathbb{R}^{c}$ and the output is concatenated with a graph level channel wise average which is expanded to populate each node of the graph. The edge attributes are the distance between nodes. The input node values always contain the node coordinates, and, for the encoder, also the data solution at the node.
In GABI: In the encoder $E^{\theta}$ we map attributed graphs to latent vectors of fixed dimensions. The decoder maps a pairing between a latent variable and a graph and populates the node attributes on the input graph. Hence the last layer in a channel wise averaging across the graph and the channel dimension is taken to be the latent dimension. The decoder $D^{\psi}$ takes as argument a graph where the node attributes are only the coordinates, hence only contains geometric information, and we expand the latent sample $\mathbf{z}$ to decode to populate all nodes with a channel dimension equal to the latent dimension.
@@ B.3.2 Additional Results
Figure 7: (a) Lose functions during training – the total loss is the sum of these. (b) histogram of the latent $p_{\mathbf{z}}^{\theta}$ across all dimensions after training.
@@ B.5.2 Additional Results
In Figure 14 we show some example geometries along with the forcing function to and resonance field to reconstruct. In Figure 15 we show 3 random draws from the prior for a given geometry to show the diversity of solutions and forcing fields.
@@ C.1 Alternative Variational Formulations
On could alternatively train a model of the form
@@
There exist other models such as UQ-VAEs (Goh et al., 2022) where the decoder is replaced by the solution operator for the assumed PDE responsible for generating the data. This of course requires us to know the governing of the problem, the boundary conditions etc – and has the same challenges and limitations as other physics-informed methodologies. Furthermore, this framework is not made to handle variable geometries.

### v3

# arXiv:2509.19929v3[stat.ML] 27 Feb 2026

@@ Abstract
Uncertainty Quantification (UQ) is paramount for inference in engineering. A common inference task is to recover full-field information of physical systems from a small number of noisy observations, a usually highly ill-posed problem. Sharing information from multiple distinct yet related physical systems can alleviate this ill-posedness. Critically, engineering systems often have complicated variable geometries prohibiting the use of standard multi-system Bayesian UQ. In this work, we introduce Geometric Autoencoders for Bayesian Inversion (GABI), a framework for learning geometry-aware generative models of physical responses that serve as highly informative geometry-conditioned priors for Bayesian inversion. Following a “learn first, observe later” paradigm, GABI distills information from large datasets of systems with varying geometries, without requiring knowledge of governing PDEs, boundary conditions, or observation processes, into a rich latent prior. At inference time, this prior is seamlessly combined with the likelihood of a specific observation process, yielding a geometry-adapted posterior distribution. Our proposed framework is architecture-agnostic. A creative use of Approximate Bayesian Computation (ABC) sampling yields an efficient implementation that utilizes modern GPU hardware. We test our method on: steady-state heat over rectangular domains; Reynolds-Averaged Navier-Stokes (RANS) flow around airfoils; Helmholtz resonance and source localization on 3D car bodies; RANS airflow over terrain. We find: the predictive accuracy to be comparable to deterministic supervised learning approaches in the restricted setting where supervised learning is applicable; UQ to be well calibrated and robust on challenging problems with complex geometries.
@@ 1 Introduction
In this work, we propose a methodology using graph-based autoencoders to learn highly informative priors as geometry-aware generative models for use in Bayesian inversion tasks. We also show how to effectively solve the Bayesian inverse problems resulting from these priors through a pushforward model. Importantly, the proposed scheme decouples the task of “learning” and the task of “inferring from observations”, resulting in train-once use-everywhere approach which is independent of the particular observation process for solving inverse problems, a trait not shared by related supervised learning approaches. We make the following contributions:
@@
    We demonstrate the scalability of the method on a large terrain flow problem through a multi-GPU implementation, showing potential for training GABI foundation models.
@@ 1.1 Setup
  • In the application stage, we use this learned prior to solve inverse problems for data coming from new physical systems not in the dataset $\mathcal{D}$. That is, given sparse and noisy observations $\mathbf{y}_{o}$ of a solution $\mathbf{u}_{o}$ of a physical system with geometry $\mathsf{M}_{o}$, we use our previously learned prior to solve the Bayesian inverse problem for the unknown full-field $\mathbf{u}_{o}$. We recover a distribution over the solution $\mathbf{u}_{o}$ by only observing $\mathbf{y}_{o}$.
Using the proposed methodology, one can encode prior beliefs through geometry-dependent full-field solution datasets. Practitioners can then construct highly-informative priors from data across varying physical geometries. This prior is independent of the observation process for any particular inference task, hence this is a learn-once apply-everywhere approach.
@@ 1.2.2 Forward and Inverse Problems on Graphs
Graphs are handled in machine learning frameworks through geometric deep learning (Bronstein et al., 2021; Wu et al., 2022; Kipf, 2016). Several important works consider inverse problems where the quantity inferred lives of a graph description of the geometry. In Garcia Trillos & Sanz-Alonso (2018) the authors prove the convergence of the graph posterior to the continuum posterior subject to some conditions. In Povala et al. (2022) the adjacency structure of a mesh is leveraged to accelerate variational inference inversion. Important works consider solving the PDE forward problem through physics-informed graph machine learning(Seo & Liu, 2019; Valencia et al., 2025; Wandel et al., 2020) Used in our comparisons are linear inverse problems with Gaussian process priors, this leads to graph Gaussian process methods (Borovitskiy et al., 2021; Mostowsky et al., 2024). We find it important to mention works in the physics-informed literature where (assuming complete or partial knowledge on PDEs, boundary conditions, forcing etc.) one can use physics residuals to learn forward and inverse maps for PDEs (Raissi et al., 2019; Gao et al., 2022; Vadeboncoeur et al., 2023; Xu et al., 2023; Rixner & Koutsourelakis, 2021). However, in this work, we focus solely on methods which do not assume knowledge of the governing PDE, boundary conditions, or forcing and are entirely data-driven in their approach.
@@ 1.3 Notation
We denote by $\mathbb{P}_{z}$ a probability measure on the measurable space $(\mathcal{Z},\mathcal{S})$, and by $p_{z}$ the associated density with respect to Lebesgue measure when it exists. Empirical measures are a mixture of Dirac measures. That is, for $D=\{x_{i}\}_{i=1}^{M}$, we have $\widehat{\mathbb{E}}_{x\in D}[\delta_{x}]=\frac{1}{M}\sum_{i=1}^{M}\delta_{x_{i}}$ where $\delta_{x_{i}}(A)=1\text{ if }x_{i}\in A,\text{ and }\delta_{x_{i}}(A)=0,\text{ if }x_{i}\notin A.$ We will sometimes refer to such probability measures as the “densities” $p,q$ to lighten notation. Superscripts in parentheses denote realizations of random variables. For a function $g:\mathcal{Z}\rightarrow\mathcal{U}$, the pushforward $g_{\#}p_{z}$ means $g_{\#}\mathbb{P}_{z}(A)=\mathbb{P}_{z}(g^{-1}(A))$ for all measurable sets $A$. We note that a sample $u^{(n)}\sim g_{\#}p_{z}$ is obtained as $u^{(n)}=g(z^{(n)}),z^{(n)}\sim p_{z}$.
@@ 2.1 Posterior from a Pushforward Prior
Stated simply, in Bayesian inference, when a prior on $\mathcal{U}$ is specified as the pushforward under $g$ (Subsection 1.3) of a latent prior on a space $\mathcal{Z}$, two distributions are equivalent: i) the Bayesian posterior on $\mathcal{U}$, and ii) the pushforward under $g$ of the posterior associated to the latent prior on $\mathcal{Z}$. This result will be explicitly related to the geometric inference setting in Theorem 2.2, showing that with a latent prior constructed via a geometry-aware encoder/decoder, we can solve inverse problems on $\mathcal{U}$ with a single common latent representation relevant to any number of in-distribution geometries. All proofs can be found in Appendix A.
@@ 3 Implementation
Neural Networks: We make use of meshes and graph neural networks to describe and manipulate geometric information. The proposed methodology is agnostic to any particular architecture, however we need the following characteristics: Geometry-aware architecture and data-structure; Mapping to and from fixed-dimensional vectors; Non-locality. We expand on these points in B. For the heat, airfoil, and car problems we make use of a Graph Convolutional Network with interleaved nonlocal averaging layers (Lanthaler et al., 2024). For the terrain flow problem we use a GEN (Li et al., 2020) GNN architecture. We detail these architectures in Appendix B.1. We also test a Transformer on the heat problem in Appendix B.3.3.
@@ 4.2 Airfoils – Flow Field
In this section, we test the proposed methodology on the full field reconstruction or airflow around an airfoil from a dataset of Reynolds averaged Navier-Stokes simulations. We use the setup described in Duthé et al. (2025). Here, the geometric autoencoder is trained on the pressure and 2D velocity field of 1k airfoil geometries with various inflow conditions. Once trained, the observation vector $\mathbf{y}_{o}$ is taken to be noisy pressure observations at a small number of nodes on the surface of the airfoil, emulating a physical scenario when such pressure sensors may be attached to a wing/blade. The Bayesian inversion task is to reconstruct the pressure and velocity field around the airfoil, a task far too ill-posed without a highly informative prior. To compile the results in Table 3 we test GABI-ABC against the Direct Map with the same architecture type – we randomize the number of observation locations to be drawn between 5 and 50 observations with probability inversely proportional to the number of observation locations. For the supervised Direct Map approach the distribution of the number of observation must be known at train time, this is not the case for GABI whose training is independent of the observation process. Figure 3 shows a select number of results. For a discussion on out-of-distribution inference we refer the reader to Appendix B.4.2.
@@ 4.3 Car Body – Acoustic Vibration and Source Localization
Figure 4: Ground truth, inferred mean, stddev., and error for amplitude $u$, and forcing $f$. In magenta are the observation locations.
@@ 5 Conclusion
We propose a methodology for learning highly informative priors for solving Bayesian inverse problems across varying geometries. The scheme is realized through a geometric autoencoder which learns informative latent representations allowing for sharing of information across geometries. Once trained this foundation model can be used with any observation process to perform Bayesian full field reconstruction over new geometries. A creative use of ABC sampling leads to an efficient GPU implementation. The method achieves similar predictive accuracy to supervised direct map methods, all while outputting a full Bayesian posterior over the fields of interest and being far more flexible due to the model being independent of the observation process. The method is tested on a wide range of physical setups and geometries.
@@ Appendix A Proofs
This follows directly from Lemma 2.1 and Theorem 2.2 by identifying $z$ with $(\mathbf{z},\sigma_{o})$, $\mathbb{P}_{z}$ with $(\mathbb{P}_{\mathbf{z}}\otimes\mathbb{P}_{\sigma_{o}})$, $g$ with $(D^{\psi}_{o}\otimes\mathrm{Id}.)$, changing the potential function to $\Phi(\mathbf{u}_{o},\sigma_{o};\mathbf{y}_{o})=\frac{1}{2\sigma_{o}^{2}}\|\mathbf{y}_{o}-\mathbf{H}_{o}\bf{\mathbf{u}}_{o}\|^{2}_{2}$, and finally noting $\mathrm{Id}._{\#}\mathbb{P}_{\sigma_{o}}=\mathbb{P}_{\sigma_{o}}$. ∎
@@ Appendix B Additional Numerical Results and Details
The proposed methodology is architecture-agnostic, however for an architecture to be appropriate for the task we require a few particular qualities:
  • Geometry-aware architecture and data-structure. As the main premise of this work is to share information about physical fields which depend on the geometry of the problem to which they relate, it is essential that the data-structure reflect this geometric information, that the neural network architecture be effective at processing it, and that both naturally accommodate inputs of varying size and topology.
  • Mapping to and from fixed-dimensional vectors. This framework relies on learning latent representations between solution to PDEs and the geometries on which these are defined. Hence, one needs to map variable-sized geometric inputs to and from fixed-dimensional latent spaces.
@@ B.1 Graph Convolutional Network
We use graph convolutional layers (Kipf, 2016) in the examples heat on rectangles, RANS around airfoil, Helmholtz on car. To induce non-locality of mappings across channels, $c$, at each layer we set $\texttt{GCN}_{l}:\mathsf{M}_{n}\times\mathbb{R}^{2c}\rightarrow\mathsf{M}_{n}\times\mathbb{R}^{c}$ and the output is concatenated with a graph level channel-wise average which is expanded to populate each node of the graph. The edge attributes are the distance between nodes. The input node values always contain the node coordinates, and, for the encoder, also the data solution at the node.
In GABI: In the encoder $E^{\theta}$ we map attributed graphs to latent vectors of fixed dimensions. The decoder maps a pairing between a latent variable and a graph and populates the node attributes on the input graph. Hence the last layer in a channel-wise averaging across the graph and the channel dimension is taken to be the latent dimension. The decoder $D^{\psi}$ takes as argument a graph where the node attributes are only the coordinates, hence only contains geometric information, and we expand the latent sample $\mathbf{z}$ to decode to populate all nodes with a channel dimension equal to the latent dimension.
@@ B.3.2 Additional Results
Figure 7: (a) Losse functions during training – the total loss is the sum of these. (b) histogram of the latent $p_{\mathbf{z}}^{\theta}$ across all dimensions after training.
@@ B.5.2 Additional Results
In Figure 14 we show some example geometries along with the forcing function and resonance field to reconstruct. In Figure 15 we show 3 random draws from the prior for a given geometry to show the diversity of solutions and forcing fields.
@@ C.1 Alternative Variational Formulations
One could alternatively train a model of the form
@@
There exist other models such as UQ-VAEs (Goh et al., 2022) where the decoder is replaced by the solution operator for the assumed PDE responsible for generating the data. This of course requires us to know the governing equations of the problem, the boundary conditions etc – and has the same challenges and limitations as other physics-informed methodologies. Furthermore, this framework is not made to handle variable geometries.